OpenStudy (anonymous):

what is the equation of the line (-2,-5) and has a slope of -4

OpenStudy (abb0t):

\(\sf \color{red}{y-y_o=m(x-x_o)}\) y\(_o\) = y point x\(_o\) = x ponit m = slope

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-2\quad ,&-5) \end{array} \\\quad \\ slope = m= -4 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

thank you.